if a man who has a mass of 75kg climbs up a mast 32m, calculate the increase in his potential energy in reaching the top. express your answer in kJ to two sig fig and take the acceleration due to gravity to be g=9.8m s^-2
1
Expert's answer
2015-09-01T01:37:56-0400
Given:
m=75 kg
delta_x=32 m
g = 9.8 m/s
deltaE-?

Solution:
deltaE=m*g*delta_x=75*32*9.8=23520 J = 23.52 kJ = 24kJ
Answer is 24 kJ